OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region under the graph of the function f(x)=x^(12/13) about the x-axis over the interval [1,3]. need some help

OpenStudy (anonymous):

V= pi integral y^2 dx

OpenStudy (anonymous):

\[ V= \pi \int\limits_{1}^{3} (x^{\frac{12}{13}})^2 dx \]

OpenStudy (anonymous):

then index laws when you raise a power to another power, you multiply

OpenStudy (anonymous):

\[V= \pi \int\limits_{1}^{3} x^{\frac{24}{13}} dx \]

OpenStudy (anonymous):

then simple

OpenStudy (anonymous):

its x ^(13/12)